The length of one leg of a right triangle is 5cm. The hypotenuse is 1 cm more than the other leg. Find the length of the hypotenuse.
Answer by edjones(8007)

a^2+b^2=c^2
b=5
c=a+1
5^2+a^2=(a+1)^2
5^2+a^2=a^2+2a+1
25=2a+1
2a=24
a=12
c=13 cm. length of the hypotenuse.
.
Ed